OpenStudy (anonymous):

please help me solve this. (36a^5/3)^1/2

OpenStudy (lgbasallote):

\(\Large \sqrt{\frac{36a^5}{3}}\) 36 can be taken out of the square root since it's a perfect square.. \(\Large 5\sqrt{\frac{a^5}{3}}\) a can be taken out too... \(\Large 5a^2 \sqrt{\frac{a}{3}}\) we can rationalize this which will give us \(\Large \frac{5a^2 \sqrt{3a}}{3}\)

OpenStudy (radar):

\[\sqrt{36^{5/3}}\] Is that the problem?

OpenStudy (lgbasallote):

seems the asker is not here...

OpenStudy (radar):

Sometimes the challenge is figuring out what the problem is lol.

OpenStudy (lgbasallote):

yes..that is a hard problem -__-

OpenStudy (anonymous):

thanks so the A power to 5/3 became A power to 5 over 3

OpenStudy (radar):

is the \[a ^{5/3}\]

OpenStudy (anonymous):

yes

OpenStudy (radar):

or is it \[a ^{3}\over 5\]

OpenStudy (lgbasallote):

my answer is wrong then lol..i thought it was the latter

OpenStudy (radar):

O.K. now it is clear.

OpenStudy (anonymous):

yeah a little bit

OpenStudy (lgbasallote):

yeah..

OpenStudy (lgbasallote):

so should you or should i? lol

OpenStudy (radar):

You do good work. Have at it, I am just an interested bystander.

OpenStudy (anonymous):

\[(36a ^{5/3})^{1/2}\] there it is in real format

OpenStudy (lgbasallote):

the way i do this is to break it down... \(\Huge (36)^\frac{1}{2} (a^\frac{5}{3})^\frac{1}{2}\) \(\Large (36)^\frac{1}{2}\) is pretty much solvable it is equal to \(\large \sqrt{36} = 6\) now \(\huge (a^\frac{5}{3})^\frac{1}{2}\) it is solveable by rule of exponents...\(\huge (a)^{(\frac{5}{3})(\frac{1}{2})}\) that is equal to...\(\huge (a)^{(\frac{5}{6})}\) combine those... \(\huge 6a^{(\frac{5}{6})}\) got it?

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

got it

OpenStudy (lgbasallote):

that was a lot of latexing lol

OpenStudy (radar):

I believe it was a good answer, well presented.

OpenStudy (lgbasallote):

\(\large \color{red}{t} \color{green}{h} \color{blue}{a} \color{orange}{n} \color{pink}{k} \color{yellow}{s}\)

OpenStudy (anonymous):

\[^{3\sqrt{ab ^{2}}}/^{3\sqrt{4a ^{2}}b}\] can u also help me rationalize the denominator in this one , please?

OpenStudy (radar):

Too small for these 73 year old eyes, and is that b in the denominator within the radical or is it out to the right ?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

within the radical

OpenStudy (radar):

\[\sqrt[3]{ab ^{2}}\over \sqrt[3]{4a ^{2}b}\] is that it or are the 3s to the left of the radical?

OpenStudy (anonymous):

yeah, that's it

OpenStudy (anonymous):

i need ur help to rationalize the denominator...

OpenStudy (radar):

To just rationalize the denominator for this fraction, multiply both the numerator and the denominator \[\sqrt[3]{4a ^{2}b}\] Since you have operated the same on the numerator and the denominator, you have not changed the value of the fraction, just the appearance.

OpenStudy (anonymous):

alright, thank you

OpenStudy (radar):

The final result of the fraction would appear similar to this: \[ab \sqrt[3]{4}\over4a ^{2}b\] which simplifies to: \[\sqrt[3]{4}\over4a\] I think.

OpenStudy (radar):

Good luck with these. It is time for bed here in Missouri.

OpenStudy (anonymous):

thanks again